Feb 01,2000 day shift


Subject: Feb 01,2000 day shift
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:49:02 CST


Shift Summary for day shift, 02/01/00

On Shift:Mike Arenton, Ricardo Barbosa + Rick K., Bob H
and others

Hours of beam delivered at all intensities: 8
Hours of beam delivered at nominal intensity: 8
Hours of KAMI data: 8

Accounting of lost time: access

===============================================================

Runs ended this shift:

Run 15723 13 hour minbias run ended
run 15725 muon run with 5 times higher gains on UCAL
run 15726 KAMI_HADRON with high gain and lower threshold on UCAL
run 15729 KAMI_HADRON normal gain, BAG in front of UCAL
run 15732 KAMI_HADRON with BAG and various absorbers

===============================================================

Tapes brought to FCC this shift: none

Accesses during shift:

move BAG into EAST beam in front of UCAL.- 0.5 hour

===============================================================

Description of Shift:

run spill time comment
----------------------------------------------------
15723 08:35 end run after 13 hours.
                   go to muon mode, took sereral short runs
                   to adjust UCAL gains up by a factor of 5.
                   and lower analog sum threshold to 30 mV.
15727 09:30 take KAMI_HADRON data with higher UCAL gains.
                   events are completely dominated by lower energy
                   junk, same with 20 inch Be absorber.
            10:15 go back to original UCAL gains.
            10:30 Bob H. and Anatoli make access to remove UCAL16
                   and move BAG into east beam in front of UCAL.
15732 11:55 run KAMI_HADRON with various absorbers.
15735 12:45 start long minbias run with 20 inch Be absorber.

   Plan is to continue this minbias run until tomorrow morning at
about 6:30 when people will enter NM2 to move magnets so as to
go to 15 mrad angle.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:49:10 CST